If the lines with equations 10x + cy - 8 = 0 and 5x - y-6 = 0 are perpendicular, find the value of c.
Answer by lwsshak3(11628)

If the lines with equations 10x + cy - 8 = 0 and 5x - y-6 = 0 are perpendicular, find the value of c.
***
rewrite given equations in standard form: y=mx+b, m=slope, b=x-intercept
10x + cy - 8 = 0
cy=-10x+8
y=-(10/c)x+8/c
..
5x - y-6 = 0
y=5x-6
..
slopes of perpendicular lines are negative reciprocals of each other
c/10=5
c=50